OpenStudy (anonymous):

Find the derivative of the inverse of the following functions at the specified point on the graph of the inverse function: f(x) = x2 + 1 for x is greater than or equal to 0. (5, 2)

myininaya (myininaya):

So we need to find the derivative of the inverse. That means we first need to find the inverse of the function that if y=x^2+1 ,x>=0

myininaya (myininaya):

do you know how to solve y=x^2+1 where x>=0 for x?

OpenStudy (anonymous):

f^-1 = sqrt(x - 1)

myininaya (myininaya):

if you don't want to do this way there is actually another way

myininaya (myininaya):

I will show you when we are done

myininaya (myininaya):

ok and then find the derivative of that by using chaining rule

myininaya (myininaya):

chain rule*

myininaya (myininaya):

if you don't know chain rule yet we should do it the other way

myininaya (myininaya):

\[\frac{d}{dx} f^{-1}(x)=\frac{1}{f'(f^{-1}(x))}\]

myininaya (myininaya):

\[\frac{d}{dx} f^{-1}(x)|_{x=5}=\frac{1}{f'(f^{-1}(5))}\]

myininaya (myininaya):

f inverse of 5 is given

myininaya (myininaya):

by that ordered pair (5,2)

OpenStudy (anonymous):

so 2?

myininaya (myininaya):

\[\frac{d}{dx} f^{-1}(x)|_{x=5}=\frac{1}{f'(f^{-1}(5))} =\frac{1}{f'(2)}\]

myininaya (myininaya):

so instead of finding the derivative of f inverse we could find the derivative of f

myininaya (myininaya):

but either way will work for this one

myininaya (myininaya):

i will show you the other way when we are done with this way

OpenStudy (anonymous):

derivative of 2 is 0.

myininaya (myininaya):

no no ...

myininaya (myininaya):

to find f'(2) you need to first find the function that is f'

myininaya (myininaya):

so f'(x)=?

OpenStudy (anonymous):

x^2 + 1 so derivative is x

myininaya (myininaya):

2x right?

OpenStudy (anonymous):

2x, yup

myininaya (myininaya):

and since f'(x)=2x then f'(2)=2(2)=?

OpenStudy (anonymous):

4

myininaya (myininaya):

\[\frac{d}{dx} f^{-1}(x)|_{x=5}=\frac{1}{f'(f^{-1}(5))} =\frac{1}{f'(2)} =\frac{1}{4}\]

myininaya (myininaya):

ok we can do it the other way too

myininaya (myininaya):

\[f^{-1}(x)=\sqrt{x-1}\]

myininaya (myininaya):

\[\frac{d}{dx}(f^{-1}(x))=\frac{1}{2\sqrt{x-1}}\]

myininaya (myininaya):

so\[\frac{d}{dx}(f^{-1}(x))|_{x=5}=\frac{1}{2\sqrt{x-1}} |_{x=5}=\frac{1}{2\sqrt{5-1}}=\frac{1}{2\sqrt{4}}=\frac{1}{2(2)}=\frac{1}{4}\]